OpenStudy (anonymous):

g'(x)= xe^-(1/2 x^2) find increasing and decreasing intervals

OpenStudy (anonymous):

\[g'(x)=\frac{x}{e^{1/2~x^2}}\] \(g'(x)\) is 0 when \(x=0\). Check the sign on \(g'(x)\) for values of \(x\) in the intervals \((-\infty,0)\) and \((0,\infty)\).